# The Case For Black Hole Remnants: A Review

### v1

@@ II How to Obtain Remnants
There are of course other routes that can possibly lead to black hole remnants, but more studies are required to check these claims. For example, many quantum gravity theories predict that at sufficiently high energy the effective dimension of spacetime is 2 instead 4 (see [48] for a review; also see [49] and the references therein). In [50] it was proposed that if the early 2-dimensional universe is governed by a dilatonic gravity theory, a black hole can exist. However, as the Universe expanded and cooled down, the spacetime transits to 3 dimensional (before it becomes 4), in 3-dimensions however, if there is no negative cosmological constant, then there can be no black hole solution. Thus, the black holes from the 2-dimensional epoch ceases its evaporation and becomes “stuck” as remnants. This picture is however, somewhat problematic. What is the geometry of a black hole remnant once it crossed the dimensional transitions? Most remnants are themselves black holes, so if no black hole solution is allowed, what are those remnants, which in any case, have two dimensions less than the surrounding space? Also, most quantum gravity results indicate that spacetime dimension does not change, but rather the spectral dimension or thermal dimension does (i.e., spacetime is only effectively 2-dimensional at high energy scale). In this case there is no problem with black holes with effective dimension being 3. While the scenario above concerns the dimensional transition of the entire Universe, we can also consider the dimensional transition of the black hole itself. As the black hole evaporates and gets hotter, eventually its spectral dimension reduces to 2. It is shown in [49] that provided that the solution still only has 1 horizon, the lifetime is only slightly longer than the standard $M^{3}$, namely $M^{3}+cM$ term, where $c$ is a constant. Therefore it seems that the spectral dimension route does not lead to remnants. However, in the case of thermal dimension [51, 52], which arises from the modification of the dispersion relation, whether remnants occur or not likely depend on the dispersion relation. Also related to dispersion relation is the idea of “gravity’s rainbow” [53], which does lead to black hole remnants [54].
@@ III Why Traditional Criticisms Against Remnants are Not Convincing
There are two major issues concerning remnants, which led to their viability being questioned by the mainstream high energy physics community. The two problems are the species problem [^7] Not to be confused with another “species problem” [55, 56] (though they may be related): if Bekenstein-Hawking entropy is entanglement entropy, it should depend on the species of particles/fields available, so why is it only proportional to the area? and the entropy problem, which are closely related. We begin with the species problem. Consider again the no-hair theorem of black holes. Once a black hole is formed, there is no way to tell from the outside what was the stuff that collapsed to form the black hole in the first place. The possibility seems to be endless, each of these different internal state gives rise to a distinct “species”. Now, consider the probability of pair-producing remnants from quantum fluctuation of the vacuum or in a high energy experiment such as a collider. For each species, whose mass is about the Planck mass, the pair-production probability is tiny. However, since there are infinitely many species, the total production rate seems to be unbounded. This therefore seems to predict that tiny black holes should be copiously produced from the vacuum all the time, as long as the available energy exceeds twice the remnant mass (given that they are Planckian, this is a relatively low energy requirement). The fact that we do not observe them therefore seemingly rules out black hole remnants. This criticism is, however over-simplistic, as we do not know if the number of internal states is actually infinite. In fact, Giddings had argued that it is finite [57, 58]. Even if there is indeed an infinitely many possible internal states, the usual physics of particle pair-production may be quite misleading when applied to remnants. Even though they may appear particle-like with mass the order of the Planck mass, they still have a vast interior structure. This is a result that follows from general relativity, not any exotic physics.
@@
For any star of radius $r$, given that it is spherical, we can deduce that its volume has to be fixed, given by $\frac{4}{3}\pi r^{3}$. However, a black hole is a highly curved spacetime manifold; its interior is nontrivial. This should be well-known but somehow the literature is still confusticated by statements about the “central singularity” of a Schwarzschild black hole, as if $r=0$ is at the geometric center of a sphere [^8] The “$r=\text{const.}$.” surfaces are sometimes referred to as “T-spheres” of the “T-region”, especially in older texts [59] to emphasize the distinction.. What happens in the Schwarzschild case is that the singularity is spacelike, an event in the future of the in-falling observer. The interior spacetime is best viewed as a special case of Kantowski-Sachs cosmological spacetime, which is anisotropic [60, 61]. In particular, inside the horizon, a line constant of constant $r,\theta,\phi$ has a proper length that goes to infinity as the singularity is approached, hence “there is no dearth of space inside” of black hole, in the words of Ted Jacobson [62].
@@
Another subtlety concerns the notion of entropy itself being observer-dependent (which is not really a surprise because it is well-known that temperature is also observer-dependent – the Hawking temperature being valid for asymptotic observers). In particular, consider a wave packet of fixed energy in flat spacetime. We can increase the entropy without adding more energy by introducing more particle species. Suppose there are some huge number $n$ of free scalar fields. Then the entropy grows like $\log n$. If we go to the Rindler wedge instead, then the von Neumann entropy of the mixed state, though increases as $\log n$, is bounded above by $E/T$ [70] (note that this is essentially the Bekenstein bound, up to a constant prefactor). This at least suggests that we should be careful with whose entropy are we talking about. Perhaps the interior of a remnant can have vastly more entropy than its surface bound, without being in tension with the fact that an asymptotic observer only sees the surface entropy. In addition, even the interpretation of the Bekenstein bound requires some care; in [71] Hayden and Wang argued that unlike classical bits and qubits, zero-bits and their associated information processing capability are generally not constrained by the Bekenstein bound.
@@ IV Discussion: Prospects for Black Hole Remnants
As we have seen, there are various ways one could obtain black hole (or white hole) remnants. It should be emphasized that the GUP approach, though relatively simple, still requires more careful investigations [27, 28, 29]. After all, there are many versions of GUP in the literature, which should we trust? Even for a fixed GUP, it is not clear what should the approach be for black hole thermodynamics. For example, should we modify only the Hawking temperature and the entropy, or also modify the metric tensor? If the latter, how do we obtain the metric? No unique, well accepted, answers are available to these questions in the literature, and as such the conclusions obtained about remnants are also subject to a certain degree of skepticism. The possibility that black holes can quantum tunnel into a white hole remnant, obtained mostly from loop quantum gravity approach, seems more promising.
@@
In principle remnants can still “resolve” the information paradox, but one has to be careful of some subtle aspects. For one, whether there is a remnant or not is quite independent from whether Hawking radiation encodes information. If information is still carried out via Hawking emission, then the firewall paradox remains because Page time sets in much earlier than the remnant stage [1] if the remnant is Planckian. However, the requirement that information be carried out by Hawking radiation rests upon the assumption of the central dogma, which as we have discussed, need not hold. Therefore, Hawking radiation can be information free like initially thought, with the information somehow contained within the large interior of the remnant. If the remnant slowly emits diffuse particles and completely evaporate in a far longer time scale (at least $M^{4}$), then eventually all the information is released back to the exterior universe and unitarity is preserved. The large interior allows these particles to be soft particles that carry little energy, which also explains why remnants are not copiously produced in everyday high energy experiments – any mini-black hole we can produce in future laboratory is a completely different object from a remnant. Much of the discussions here are not new, unfortunately the current state remains more or less the same, with many colleagues still unaware of the counter-arguments against the species and entropy problems.
@@
The interior volume of black holes did however, gained some attention in the high energy community due to it possibly being a holographic dual to complexity (the so-called “Complexity-Volume conjecture” or just CV-conjecture [72, 73, 74, 75, 76, 77, 78]), yet how this might relate to a remnant is still unclear. In [79], Susskind proposed the idea that black hole complexity cannot grow without bound. Instead, after some time that scales as $\exp{(S)}$, where $S$ is the Bekenstein-Hawking entropy [^9] This time scale, exponential in the entropy, is also the time scale that is required for one to decode the information contained in Hawking radiation [80], assuming the latter contains information. , the quantum state of the black hole now also contains a significant white hole component. This is the “gray hole” phase, during which the complexity fluctuates around a constant value. After a time of $t\sim\exp[{\exp{(S)}}]$ the “gray hole” becomes a white hole, during which the complexity decreases. Note that this is very similar to the black-to-white hole transition we have discussed, but the time scale involved in extremely different. The possibility that even astrophysical black holes are “gray”, i.e., in the quantum superposition of black and white hole states has also been discussed in [81], in which the authors also pointed out it is possible to arrange the “black-to-white transition” to have zero action – so that it will not be subject to destructive interference in the Feynman path integral.
@@
We should also not rule out the possibility that information escapes into “another universe”, interpreted broadly. For example, many of the regular black hole solutions have de-Sitter core or other geometries (for example, an interior which asymptotes to $dS_{2}\times S^{2}$ in the future is obtained in [82]), which essentially is like another universe whose spacetime dimension becomes effectively 2-dimensional (the size of the $S^{2}$ is fixed). Even under the usual Hawking evaporation without remnant, we already mentioned that the interior volume continues to grow. Therefore, if the black hole completely evaporates away, the interior spacetime must pinch off its parent (our Universe) and subsequently becomes a new baby universe, taking away all the information with it (assuming that Hawking radiation is information free) [83]. The idea that black holes lead to a new universe and thus avoiding the information loss problem is of course, not new either, see, e.g., [84]. The pinching-off of a baby universe is sometimes referred to as the black hole “emitting” a baby universe. In [85], based on a calculation carried out in Jackiw-Teitelboim (JT) gravity, it was argued that a very old black hole can tunnel to a white hole or even a firewall by emitting a large baby universe. However, the probability for tunneling to a white hole is proportional to $t^{2}\exp{(-2S)}$, where t is the age of the black hole, which again differs from the loop quantum gravity tunneling probability of $\exp{(-S)}$ [47] (granted that the gravity theories are different). In addition, the loop quantum gravity tunneling does not “emit” away the large interior; the interior just slowly shrinks throughout the white hole phase until the remnant completely disappears via diffuse emission.
@@
It is also interesting to compare the remnant picture with that of the usual remnant-free end stage of Hawking evaporation. Note that as a Schwarzschild black hole evaporates, it gets smaller, which in turn means that the curvature (say, the Kretschmann scalar) at and near its horizon is becoming larger. In a truly quantum picture $M$ cannot be continuously changed (although semi-classically we model the evaporation with a differential equation of $dM/dt$), so at the final stage the Planckian black hole is expected to dissolve away into a collection of particles and we are left with a flat spacetime. Still, this is only an expectation. How do we know that the large spacetime curvature does not “tear” spacetime enough to lead to a naked singularity? Certainly such a possibility has been hinted at in [86, 87]. Including higher curvature terms (quadratic gravity) also suggests such a possibility [88], though it is hard to say what the full quantum gravity result would be. We normally expect that the end state is flat spacetime because it is the smooth limit of $M\to 0$, but it is not a smooth limit for the curvature. Of course, naked singularity as the end state will be bad, and by the (weak) cosmic censorship conjecture it is natural to think that this does not happen. Still, this does not completely avoid the problem, as we will discuss below.
@@
In the remnant picture, if the lifetime is finite, the same thing happens towards the end. We do not know for sure whether quantum gravity can indeed “resolve” singularity completely [^10] Note that GUP may or may not remove the classical singularity. The remnant obtained in [25] still considers the same classical form of Schwarzschild geometry, so the singularity is untouched (here the “minimal length” is reflected in the size of the remnant, not in “smearing out” the singularity. For a review about minimal length, see [89].). Though in other models, one could also employ GUP to obtain regular black holes [90, 91]., or replace it with a milder, quantum version (see [92, 93] for my related reviews). In fact, if a remnant has more entropy than the Bekenstein-Hawking entropy, then a version of singularity theorem can be proved [94]. This means that there is still the possibility that the end state is either a singularity or something else that can only be described by quantum gravity. As Penrose wrote in [95]:
@@
Even if the lifetime of the black hole is infinite, one cannot avoid this problem as long as the mass becomes Planckian. This is because naked timelike or null singularities are bad in classical general relativity not because of the singularities themselves, but because their presence renders the theory non-predictive. Without knowing the properties of the singularities, one cannot know what influences it can have on spacetime region even far away from them. Note that this problem does not only arise when the curvature is mathematically divergent, it already arises for large enough curvatures. Usually we expect general relativity to be a good theory when the curvature scale is above the Planck scale. Thus, if black holes become Planckian in size (whether there is a Planck size remnant or not), cosmic censorship will be violated in spirit. Again, it is worth emphasizing that stable remnants (those with infinite lifetime) cannot serve as a cosmic censor, it effectively becomes a “Planckian curvature source”, as bad as a naked singularity. This is not necessarily a disaster, it just means that the end state of Hawking evaporation cannot be understood semi-classically. Remnants that stop evaporating after either the Page time or scrambling time is effectively classical and this problem never arises. One might think that this problem also does not arise for extremal black hole remnants if they asymptote to extremal state when still being macroscopic in size. However, this is not the case, as various studies have suggested that near-extremal black holes are highly quantum object and thus the result cannot be trusted [96, 97, 98, 99, 100, 101, 102]. In fact, an extremal black hole can be effectively a singular object [96, 98, 99].
@@
To conclude, no arguments against remnants hold up to close scrutiny, and thus the option that remnants can avoid information paradox remains viable. Given that Planck size remnants may also serve as a dark matter candidate (see, e.g., [103, 105, 104, 42]) — a possibility already noted by Aharonov, Casher, and Nussinov [2] — it would be interesting to further investigate their properties under different models. For example, recently there was a concern that some asymptotically extremal black hole remnants may be too fast (due to the recoil of the emitted radiation) [106] and thus too warm to be dark matter, but subsequent discussions had reassured that it is still viable [107, 108].
@@
In the complete evaporation picture, the lower bound of the primordial black hole mass is set by the Hawking evaporation rate, namely that black holes cannot be too small or otherwise they would have completely evaporated by now. This yields a lower bound of about $10^{17}$ g. With remnants, this bound has to be modified, depending on the model. For example in [47], the white hole remnant is estimated to be in the mass range of $10^{10}$ g and $10^{15}$ g, and Schwarzschild radius in the range $10^{-18}$ cm and $10^{-13}$ cm. These bounds are obtained from the requirement that the white hole remnant lifetime is longer than the Hubble time, and its progenitor black hole’s lifetime is shorter than the Hubble time, namely $M^{3}<t_{\text{Hubble}}<M^{4}$. If, however, black holes already stop evaporating after the Page time or even the scrambling time, then again we must re-check the bounds. In [109], it is shown that if the remnant forms after about half of its mass is evaporated (the “memory burden model”), then we could have light primordial black holes below $10^{10}$ g serving as a dark matter candidate. Some mass ranges of remnants can hopefully be tested in the future via gravitational wave signature by, e.g., the Einstein Telescope. One such scenario was studied in [110], in which it was pointed out that if the remnant mass is around $5\times 10^{5}$ g, then there will be a cosmological gravitational wave signal at frequencies around 100 Hz. What is the actual mass scale for a black hole transitioning into a remnant is therefore an important question that requires more studies.

### v2

@@ II How to Obtain Remnants
There are of course other routes that can possibly lead to black hole remnants, but more studies are required to check these claims. For example, many quantum gravity theories predict that at sufficiently high energy the effective dimension of spacetime is 2 instead of 4 (see [57] for a review; also see [58] and the references therein). In [59] it was proposed that if the early 2-dimensional universe is governed by a dilatonic gravity theory, a black hole can exist. However, as the universe expanded and cooled down, the spacetime transited into 3 dimensions (before it becomes 4). In 3-dimensions however, if there is no negative cosmological constant, then there can be no black hole solution [60]. Thus, the black holes from the 2-dimensional epoch ceased its evaporation and became “stuck” as remnants. This picture is however, somewhat problematic. What is the geometry of a black hole remnant once it crossed the dimensional transitions? Most remnants are themselves black holes, so if no black hole solution is allowed, what are those remnants, which in any case, have two dimensions less than the surrounding space? Also, most quantum gravity results indicate that spacetime dimension does not change, but rather the spectral dimension or thermal dimension does (i.e., spacetime is only effectively 2-dimensional at high energy scale). In this case there is no problem with black holes with effective dimension being 3. While the scenario above concerns the dimensional transition of the entire universe, we can also consider the dimensional transition of the black hole itself. As the black hole evaporates and gets hotter, eventually its spectral dimension reduces to 2. It is shown in [58] that provided that the solution still only has 1 horizon, the lifetime is only slightly longer than the standard $M^{3}$, namely $M^{3}+cM$ term, where $c$ is a constant. Therefore it seems that the spectral dimension route does not lead to remnants. However, in the case of thermal dimension [61, 62], which arises from the modification of the dispersion relation, whether remnants occur or not likely depend on the dispersion relation. Also related to dispersion relation is the idea of “gravity’s rainbow” [63], which does lead to black hole remnants [64].
@@ III Why Traditional Criticisms Against Remnants are Not Convincing
There are two major issues concerning remnants, which led to their viability being questioned by the mainstream high energy physics community. The two problems are the species problem [^12] Not to be confused with another “species problem” [65, 66] (though they may be related): if Bekenstein-Hawking entropy is entanglement entropy, it should depend on the species of particles/fields available, so why is it only proportional to the area? and the entropy problem, which are closely related. We begin with the species problem. Consider again the no-hair theorem of black holes. Once a black hole is formed, there is no way to tell from the outside what was the stuff that collapsed to form the black hole in the first place. The possibility seems to be endless, each of these different internal states gives rise to a distinct “species”. Now, consider the probability of pair-producing remnants from quantum fluctuation of the vacuum or in a high energy experiment such as a collider. For each species, whose mass is about the Planck mass, the pair-production probability is tiny. However, since there are infinitely many species, the total production rate seems to be unbounded. This therefore seems to predict that tiny black holes should be copiously produced from the vacuum all the time, as long as the available energy exceeds twice the remnant mass (given that they are Planckian, this is a relatively low energy requirement). The fact that we do not observe them therefore seemingly rules out black hole remnants. This criticism is, however, over-simplistic, as we do not know if the number of internal states is actually infinite. In fact, Giddings had argued that it is finite [67, 68]. Even if there is indeed an infinitely many possible internal states, the usual physics of particle pair-production may be quite misleading when applied to remnants. Even though they may appear particle-like with mass the order of the Planck mass, they still have a vast interior structure. The vast amount of space inside a black hole is a result that follows from standard general relativity, not any exotic physics.
@@
For any star of radius $r$, given that it is spherical, we can deduce that its volume has to be fixed, given by $\frac{4}{3}\pi r^{3}$. However, a black hole is a highly curved spacetime manifold; its interior is nontrivial. This should be well-known but somehow the literature is still confusticated by statements about the “central singularity” of a Schwarzschild black hole, as if $r=0$ is at the geometric center of a sphere [^13] The “$r=\text{const.}$.” surfaces inside the black hole are sometimes referred to as “T-spheres” of the “T-region”, especially in older texts [69] to emphasize the distinction.. What happens in the Schwarzschild case is that the singularity is spacelike, an event in the future of the in-falling observer. The interior spacetime is best viewed as a special case of Kantowski-Sachs cosmological spacetime, which is anisotropic [70, 71]. In particular, inside the horizon, a line of constant $r,\theta,\phi$ has a proper length that goes to infinity as the singularity is approached, hence “there is no dearth of space inside” of a black hole, in the words of Ted Jacobson [72].
@@
Another subtlety concerns the notion of entropy itself being observer-dependent (which is not really a surprise because it is well-known that temperature is also observer-dependent – the Hawking temperature being valid for asymptotic observers). In particular, consider a wave packet of fixed energy in flat spacetime. We can increase the entropy without adding more energy by introducing more particle species. Suppose there are some huge number $n$ of free scalar fields. Then the entropy grows like $\log n$. If we go to the Rindler wedge instead, then the von Neumann entropy of the mixed state, though increases as $\log n$, is bounded above by $E/T$ [80] (note that this is essentially the Bekenstein bound, up to a constant prefactor). This at least suggests that we should be careful about whose entropy are we talking about. Perhaps the interior of a remnant can have vastly more entropy than its surface bound, without being in tension with the fact that an asymptotic observer only sees the surface entropy. In addition, even the interpretation of the Bekenstein bound requires some care; in [81] Hayden and Wang argued that unlike classical bits and qubits, zero-bits [^14] A zero-bit is a special case of $\alpha$-bits, introduced in [82]. They correspond to the ability to perform quantum error correction on arbitrary bounded dimension subspaces, with the real number $0\leqslant\alpha\leqslant 1$ characterizing the size of the subspace. The standard quantum error correction corresponds to $\alpha=1$, while $\alpha=0$ is related to quantum identification, a form of equality testing for quantum states. In the context of black holes, if the central dogma holds, the zero-bits are revealed after the Page time [83]. The information scrambling of Hayden-Preskill protocol [39] also corresponds to the $\alpha=0$ case [83]. and their associated information processing capability are generally not constrained by the Bekenstein bound. In other words, we must understand the various subtleties of entropy before we can definitely rule out remnants.
@@ IV Discussion: Prospects for Black Hole Remnants
As we have seen, there are various ways one could obtain black hole (or white hole) remnants. It should be emphasized that the GUP approach, though relatively simple, still requires more careful investigations [35, 109, 37]. After all, there are many versions of GUP in the literature, which should we trust? Even for a fixed GUP, it is not clear what should the approach be for black hole thermodynamics. For example, should we modify only the Hawking temperature and the entropy, or also modify the metric tensor? If the latter, how do we obtain the metric? No unique, well accepted, answers are available to these questions in the literature [109], and as such the conclusions obtained about remnants are also subject to a certain degree of skepticism. The possibility that black holes can quantum tunnel into a white hole remnant, obtained mostly from loop quantum gravity approach, seems more promising.
@@
In principle remnants can still “resolve” the information paradox, but one has to be careful of some subtle aspects. For one, whether there is a remnant or not is quite independent from whether Hawking radiation encodes information. If information is still carried out via Hawking emission, then the firewall paradox remains because Page time sets in much earlier than the remnant stage [1] if the remnant is Planckian. However, the requirement that information be carried out by Hawking radiation rests upon the assumption of the central dogma, which as we have discussed, need not hold. Therefore, Hawking radiation can be information free like initially thought, with the information somehow contained within the large interior of the remnant. If the remnant slowly emits diffuse particles and completely evaporates in a far longer time scale (at least $M^{4}$), then eventually all the information is released back to the exterior universe and unitarity is preserved. The large interior allows these particles to be soft particles that carry little energy, which also explains why remnants are not copiously produced in everyday high energy experiments – any mini-black hole we can produce in future laboratory is a completely different object from a remnant. Much of the discussions here are not new, unfortunately the current state remains more or less the same, with many colleagues still unaware of the counter-arguments against the species and entropy problems.
@@
The interior volume of black holes did however, gained some attention in the high energy community due to it possibly being a holographic dual to complexity (the so-called “Complexity-Volume conjecture” or just CV-conjecture [84, 85, 86, 87, 88, 89, 90]), yet how this might relate to a remnant is still unclear. In [91], Susskind proposed the idea that black hole complexity cannot grow without bound. Instead, after some time that scales as $\exp{(S)}$, where $S$ is the Bekenstein-Hawking entropy [^15] This time scale, exponential in the entropy, is also the time scale that is required for one to decode the information contained in Hawking radiation [93], assuming the latter contains information. , the quantum state of the black hole now also contains a significant white hole component. This is the “gray hole” phase [^16] A different kind of “gray hole” is considered in [92]., during which the complexity fluctuates around a constant value. After a time of $t\sim\exp[{\exp{(S)}}]$ the “gray hole” becomes a white hole, during which the complexity decreases. Note that this is very similar to the black-to-white hole transition we have discussed, but the time scale involved is extremely different. The possibility that even astrophysical black holes are “gray”, i.e., in the quantum superposition of black and white hole states has also been discussed in [94], in which the authors also pointed out it is possible to arrange the “black-to-white transition” to have zero action – so that it will not be subject to destructive interference in the Feynman path integral.
@@
We should also not rule out the possibility that information escapes into “another universe”, interpreted broadly. For example, many of the regular black hole solutions have de-Sitter core or other geometries (for example, an interior which asymptotes to $dS_{2}\times S^{2}$ in the future is obtained in [95]), which essentially is like another universe whose spacetime dimension becomes effectively 2-dimensional (the size of the $S^{2}$ is fixed). Even under the usual Hawking evaporation without remnant, we already mentioned that the interior volume continues to grow. Therefore, if the black hole completely evaporates away, the interior spacetime must pinch off its parent (our universe) and subsequently becomes a new baby universe, taking away all the information with it (assuming that Hawking radiation is information free) [96]. The idea that black holes lead to a new universe and thus avoiding the information loss problem is of course, not new either; see, e.g., [97]. The pinching-off of a baby universe is sometimes referred to as the black hole “emitting” a baby universe. In [98], based on a calculation carried out in Jackiw-Teitelboim (JT) gravity, it was argued that a very old black hole can tunnel to a white hole or even a firewall by emitting a large baby universe. However, the probability for tunneling to a white hole is proportional to $t^{2}\exp{(-2S)}$, where t is the age of the black hole, which again differs from the loop quantum gravity tunneling probability of $\exp{(-S)}$ [55] (granted that the gravity theories are different). In addition, the loop quantum gravity tunneling does not “emit” away the large interior; the interior just slowly shrinks throughout the white hole phase until the remnant completely disappears via diffuse emission.
@@
It is also interesting to compare the remnant picture with that of the usual remnant-free end stage of Hawking evaporation. Note that as a Schwarzschild black hole evaporates, it gets smaller, which in turn means that the curvature (say, the Kretschmann scalar) at and near its horizon is becoming larger. In a truly quantum picture $M$ cannot be continuously changed (although semi-classically we model the evaporation with a differential equation of $dM/dt$), so at the final stage the Planckian black hole is expected to dissolve away into a collection of particles and we are left with a flat spacetime. Still, this is only an expectation. How do we know that the large spacetime curvature does not “tear” spacetime enough to lead to a naked singularity? Certainly such a possibility has been hinted at in [99, 100]. Including higher curvature terms (quadratic gravity) also suggests such a possibility [101], though it is hard to say what the full quantum gravity result would be. We normally expect that the end state is flat spacetime because it is the smooth limit of $M\to 0$, but it is not a smooth limit for the curvature. Of course, naked singularity as the end state will be bad, and by the (weak) cosmic censorship conjecture it is natural to think that this does not happen. Perhaps remnants also play the role of a cosmic censor. Still, this does not completely avoid the problem, as we will discuss below.
@@
In the remnant picture, if the lifetime is finite, the same thing happens towards the end. We do not know for sure whether quantum gravity can indeed “resolve” singularity completely [^17] Note that GUP may or may not remove the classical singularity. The remnant obtained in [33] still considers the same classical form of Schwarzschild geometry, so the singularity is untouched (here the “minimal length” is reflected in the size of the remnant, not in “smearing out” the singularity. For a review about minimal length, see [102].). Though in other models, one could also employ GUP to obtain regular black holes [103, 104]., or replace it with a milder, quantum version (see [105, 106] for my related reviews). In fact, if a remnant has more entropy than the Bekenstein-Hawking entropy, then a version of singularity theorem can be proved [107]. This means that it is still possible that the end state is either a singularity or something else that can only be described by quantum gravity. As Penrose wrote [108]:
@@
Even if the lifetime of the black hole is infinite, one cannot avoid this problem as long as the mass becomes Planckian. This is because naked timelike or null singularities are bad in classical general relativity not because of the singularities themselves, but because their presence renders the theory non-predictive. Without knowing the properties of the singularities, one cannot know what influences it can have on spacetime region even far away from them. Note that this problem does not only arise when the curvature is mathematically divergent, it already arises for large enough curvatures. Usually we expect general relativity to be a good theory when the curvature scale is above the Planck scale (for the subtleties, see [109]). Thus, if black holes become Planckian in size (whether there is a Planck size remnant or not), cosmic censorship will be violated in spirit. Again, it is worth emphasizing that stable remnants (those with infinite lifetime) cannot serve as a cosmic censor, it effectively becomes a “Planckian curvature source”, as bad as a naked singularity. This is not necessarily a disaster, it just means that the end state of Hawking evaporation cannot be understood semi-classically. Nevertheless, we should keep in mind that one of the criticisms against firewall is that it is essentially a naked singularity [11]. Should we prefer remnants over firewalls if they have the same problems? Perhaps the former is more palatable because it only arises for Planckian black holes, whereas the latter can occur even for large black holes.
Remnants that stop evaporating after either the Page time or scrambling time is effectively classical and this issue never arises. One might think that this problem also does not arise for extremal black hole remnants if they asymptote to extremal state when still being macroscopic in size. However, this is not the case, as various studies have suggested that near-extremal black holes are highly quantum object and thus the result cannot be trusted [110, 111, 112, 113, 114, 115, 116]. In fact, an extremal black hole can be effectively a singular object [110, 112, 113].
@@
To conclude, no argument against remnants holds up to close scrutiny, and thus the option that remnants can avoid information paradox remains viable. Given that Planck size remnants may also serve as a dark matter candidate (see, e.g., [117, 118, 119, 50]) — a possibility already noted by Aharonov, Casher, and Nussinov [2] — it would be interesting to further investigate their properties under different models. For example, recently there was a concern that some asymptotically extremal black hole remnants may be too fast (due to the recoil of the emitted radiation) [121] and thus too warm to be dark matter, but subsequent discussions had reassured that it is still viable [122, 123].
@@
In the complete evaporation picture, the lower bound of the primordial black hole mass is set by the Hawking evaporation rate, namely that black holes cannot be too small or otherwise they would have completely evaporated by now. This yields a lower bound of about $10^{17}$ g. With remnants, this bound has to be modified, depending on the model. For example in [55], the white hole remnant is estimated to be in the mass range of $10^{10}$ g and $10^{15}$ g, and Schwarzschild radius in the range $10^{-18}$ cm and $10^{-13}$ cm. These bounds are obtained from the requirement that the white hole remnant lifetime is longer than the Hubble time, and its progenitor black hole’s lifetime is shorter than the Hubble time, namely $M^{3}<t_{\text{Hubble}}<M^{4}$. If, however, black holes already stop evaporating after the Page time or even the scrambling time, then again we must re-check the bounds. In [124], it is shown that if the remnant forms after about half of its mass is evaporated (the “memory burden model”), then we could have light primordial black holes below $10^{10}$ g serving as a dark matter candidate. It has recently been pointed out that the evaporation rate for cosmological black holes cannot be treated the same way as isolated black holes in asymptotically flat spacetimes. The consequence is that the decay of primordial black holes occurs faster for larger masses but the decay rate reduces for lower mass [125]. Some mass ranges of remnants can hopefully be tested in the future via gravitational wave signature by, e.g., the Einstein Telescope. One such scenario was studied in [126], in which it was pointed out that if the remnant mass is around $5\times 10^{5}$ g, then there will be a cosmological gravitational wave signal at frequencies around 100 Hz. The aforementioned memory-burden proposal can potentially also be tested via gravitational waves [127]. What is the actual mass scale for a black hole transitioning into a remnant is therefore an important question that requires more studies.